Find the value of k so that the lines 2x – 3y = 9 and kx-9y =18 will be parallel.
Answer by stanbon(75887)

Find the value of k so that the lines 2x – 3y = 9 and kx-9y =18 will be parallel.
--------------
2x – 3y = 9
slope = 2/3
-------------------
and kx-9y =18
slope = k/9
-----
Solve:
k/9 = 2/3
k = 6
====================
Cheers,
Stan H.
====================